LITERATÚRA

[A] T. Aubin: Nonlinear analysis on manifolds, Monge-Ampère equations. Springer, Berlin - Heidelberg - New York, 1982

[C] J.B. Conway: A course in functional analysis. Springer, New York - Berlin - Heidelberg - Tokyo, 1985

[CH] R. Courant & D. Hilbert: Methods of mathematical physics. Interscience, New York, 1953, 1962

[D] M.M. Day: Normed linear spaces. Springer, Berlin - Heidelberg - New York, 1973 [DL] R. Dautray & J.L. Lions: Analyse mathématique et calcul numérique pour les sciences

et les techniques. Masson, Paris, 1985 [DS] N. Dunford & J.T. Schwartz: Linear operators. Interscience Publishers, New York -

London, 1958, 1963, 1971

[E] R.E. Edwards: Functional analysis. Holt, Rinehart & Winston, New York, 1965 [F] H. Federer: Geometric measure theory. Springer, New York, 1969

[FM] S. Fučík & J. Milota: Matematická analýza II (skriptá MFF UK Praha). SPN, Praha, 1980

[Gb] S. Goldberg: Unbounded linear operators. McGraw-Hill, New York, 1966

[Gs] J.A. Goldstein: Semigroups of linear operators and applications. Oxford University Press, Oxford, 1985

[Ha] P.R. Halmos: A Hilbert space problem book. Springer, Berlin - Heidelberg - New York, 1982

[He] D. Henry: Geometric theory of semilinear parabolic equations. Springer, Berlin - Heidelberg - New York, 1981

[HP] E. Hille & R.S. Phillips: Functional analysis and semi-groups. AMS, Providence, 1957 [HS] E. Hewitt & K. Stromberg: Real and abstract analysis. Springer, Berlin - Heidelberg

- New York, 1965

[JZ] K. John & V. Zizler: Topologické lineární prostory (skriptá MFF UK Praha). SPN, Praha, 1978

[K] G. Köthe: Topological vector spaces. Springer, New York, 1969

[Ka] T. Kato: Perturbation theory for linear operators. Springer, Berlin - Heidelberg - New York, 1980

[KA] L.V. Kantorovič & G.P. Akilov: Funkcional’nyj analiz. Nauka, Moskva, 1977 [KG] A.A. Kirillov & A.D. Gvišiani: Teoremy i zadači funkcional’novo analiza. Nauka,

Moskva, 1979

[KF] A.N. Kolmogorov & S.V. Fomin: Elementy teorii funkcij i funkcional’novo analiza. Nauka, Moskva, 1968

[KJF] A. Kufner & O. John & S. Fučík: Function spaces. Academia, Praha, 1977 [M] L. Mišík: Funkcionálna analýza. Alfa, Bratislava, 1989

[NS] A.W. Naylor & G.R. Sell: Linear operator theory in engineering and science. Holt, Rinehart & Winston, New York, 1971 (slovenský preklad Alfa, Bratislava, 1981)

[P] A. Pazy: Semigroups of linear operators and applications to partial differential equations. Springer, New York - Berlin - Heidelberg - Tokyo, 1983

[π1] J. Lukeš: Teorie míry a integrálu I (skriptá MFF UK Praha). SPN, Praha, 1977 [π2] J. Lukeš: Příklady z matematické analýzy I (skriptá MFF UK Praha). SPN, Praha,

1972

[Ri] R.D. Richtmayer: Principles of advanced mathematical physics. Springer, New York - Heidelberg - Berlin, 1978, 1981

[R1] W. Rudin: Real and complex analysis. McGraw–Hill, New York, 1974

[R2] W. Rudin: Functional analysis. McGraw-Hill, New York, 1973

[RN] F. Riesz & B. Sz.–Nagy: Functional analysis. Ungar, New York, 1955

[RS] M. Reed & B. Simon: Methods of modern mathematical physics. Academic Press, New York - London, 1972, 1975, 1978, 1979

[S1] H.H. Schaefer: Topological vector spaces. McMillan, New York, 1966

[S2] H.H. Schaefer: Banach lattices and positive operators. Springer, Berlin - Heidelberg - New York, 1974

[Sc] L. Schwartz: Méthodes mathématiques pour les sciences physiques. Hermann, Paris, 1965 (český preklad SNTL, Praha, 1972)

[T] A.E. Taylor: Introduction to functional analysis. Wiley, New York, 1967 (český preklad Academia, Praha, 1973)

[Tr] V.A. Trenogin: Funkcional’nyj analiz. Nauka, Moskva, 1980

[TPS] V.A. Trenogin & B.M. Pisarevskij & T.S. Soboleva: Zadači i upražnenija po funkcio-

nal’nomu analizu. Nauka, Moskva, 1984

[Y] K. Yosida: Functional analysis. Springer, Berlin - Göttingen - Heidelberg, 1965